A volleyball is sent over a net, having been launched at a speed 4.4 m/s at an angle 26.7 degrees above the horizontal, from the ground. What is the ball’s speed, in m/s, 2.66 seconds after being launched?
Answer #1

speed 4.4 m/s

angle 26.7 degrees

t = 2.66 secs.

horizomtal component

v= vox=4.4cos26.7 = 0.89m/s

vy = vyo-gt = [4.4sin26.7-9.8*2.66]m/s = -24.09m/s

speed = [0.892+24.092]1/2m/s = 24.11m/s